40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 8 19:54:06 2012, Manasa, Configuration, IOO, MC trans optics configured Pic1.png
    Reply  Wed Aug 8 22:12:30 2012, rana, Configuration, IOO, MC trans optics configured 
       Reply  Wed Aug 8 22:17:43 2012, Manasa, Configuration, IOO, MC trans optics configured 
          Reply  Fri Aug 10 09:54:51 2012, Manasa, Configuration, IOO, MC trans optics configured 
             Reply  Fri Aug 10 11:05:33 2012, jamie, Configuration, IOO, MC trans optics configured 
                Reply  Fri Aug 10 15:05:52 2012, Manasa, Configuration, IOO, MC trans optics configured 
                   Reply  Thu Aug 16 17:28:51 2012, Manasa, Configuration, IOO, MC trans optics configured 
                      Reply  Tue Oct 2 14:20:33 2012, Manasa, Configuration, IOO, PDA255 not working 
Message ID: 7206     Entry time: Thu Aug 16 17:28:51 2012     In reply to: 7144     Reply to this: 7462
Author: Manasa 
Type: Configuration 
Category: IOO 
Subject: MC trans optics configured 

Quote:

Quote:

Quote:

Quote:

Quote:

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

 We found a PDA255 but it doesn't seem to work. I am not sure if that is one you are mentioning...but I'll ask Steve tomorrow!

 I double checked the PDA255 found at the 40m and it is broken/bad. Also there was no success hunting PDs at Bridge. So the MC trans is still in the same configuration. Nothing has changed. I'll try doing ringdown measurements with PDA400 today.

Can you explain more what "broken/bad" means?  Is there no signal?  Is it noisy?  Glitch?  etc.

 The PD saturates the oscilloscope just by switching on the power; with no real signal at all. But Steve helped locating a PD that is not being used at the AP table. So I will check it and replace the current one if it works!

Koji opened up the PD and found that the screw connecting the PD to the pole was doing an additional job as well; connecting the power cable to the PD output in the inside. The PD is now fixed! Yippie...we have two PDA255 s at 40m now!!

ELOG V3.1.3-